OpenStudy (anonymous):

using the substitution u=lne find ∫1e(lnx)/x.dx

OpenStudy (anonymous):

think you mean \[\int\limits \frac{\ln(x)}{x}dx\]

OpenStudy (anonymous):

u=ln(x) so du = dx/ x